\providecommand{\href}[2]{#2}\begingroup\raggedright\begin{thebibliography}{10}

\bibitem{Spiridonov:2010em}
V.~P. Spiridonov, ``{Elliptic beta integrals and solvable models of statistical
  mechanics},'' {\em Contemp. Math.} {\bfseries 563} (2012) 181--211,
\href{http://arxiv.org/abs/1011.3798}{{\ttfamily arXiv:1011.3798 [hep-th]}}.
%%CITATION = ARXIV:1011.3798;%%.

\bibitem{Yamazaki:2012cp}
M.~Yamazaki, ``{Quivers, YBE and 3-manifolds},''
  \href{http://dx.doi.org/10.1007/JHEP05(2012)147}{{\em JHEP} {\bfseries 05}
  (2012) 147},
\href{http://arxiv.org/abs/1203.5784}{{\ttfamily arXiv:1203.5784 [hep-th]}}.
%%CITATION = ARXIV:1203.5784;%%.

\bibitem{Gahramanov:2015cva}
I.~Gahramanov and V.~P. Spiridonov, ``{The star-triangle relation and 3d
  superconformal indices},''
  \href{http://dx.doi.org/10.1007/JHEP08(2015)040}{{\em JHEP} {\bfseries 08}
  (2015) 040},
\href{http://arxiv.org/abs/1505.00765}{{\ttfamily arXiv:1505.00765 [hep-th]}}.
%%CITATION = ARXIV:1505.00765;%%.

\bibitem{Gahramanov:2016ilb}
I.~Gahramanov and A.~P. Kels, ``{The star-triangle relation, lens partition
  function, and hypergeometric sum/integrals},''
  \href{http://dx.doi.org/10.1007/JHEP02(2017)040}{{\em JHEP} {\bfseries 02}
  (2017) 040},
\href{http://arxiv.org/abs/1610.09229}{{\ttfamily arXiv:1610.09229 [math-ph]}}.
%%CITATION = ARXIV:1610.09229;%%.

\bibitem{Gahramanov:2017ysd}
I.~Gahramanov and S.~Jafarzade, ``{Integrable lattice spin models from
  supersymmetric dualities},''
  \href{http://dx.doi.org/10.1134/S1547477118060079}{{\em Phys. Part. Nucl.
  Lett.} {\bfseries 15} no.~6, (2018) 650--667},
\href{http://arxiv.org/abs/1712.09651}{{\ttfamily arXiv:1712.09651 [math-ph]}}.
%%CITATION = ARXIV:1712.09651;%%.

\bibitem{Kels:2017vbc}
A.~P. Kels and M.~Yamazaki, ``{Lens elliptic gamma function solution of the
  Yang–Baxter equation at roots of unity},''
  \href{http://dx.doi.org/10.1088/1742-5468/aaa8fd}{{\em J. Stat. Mech.}
  {\bfseries 1802} no.~2, (2018) 023108},
\href{http://arxiv.org/abs/1709.07148}{{\ttfamily arXiv:1709.07148 [math-ph]}}.
%%CITATION = ARXIV:1709.07148;%%.

\bibitem{Yamazaki:2018xbx}
M.~Yamazaki, ``{Integrability as Duality: the Gauge/YBE Correspondence},''
\href{http://arxiv.org/abs/1808.04374}{{\ttfamily arXiv:1808.04374 [hep-th]}}.
%%CITATION = ARXIV:1808.04374;%%.

\bibitem{Bozkurt:2020gyy}
D.~N. Bozkurt, I.~Gahramanov, and M.~Mullahasanoglu, ``{Lens partition
  function, pentagon identity, and star-triangle relation},''
  \href{http://dx.doi.org/10.1103/PhysRevD.103.126013}{{\em Phys. Rev. D}
  {\bfseries 103} no.~12, (2021) 126013},
  \href{http://arxiv.org/abs/2009.14198}{{\ttfamily arXiv:2009.14198
  [hep-th]}}.

\bibitem{de-la-Cruz-Moreno:2020xop}
J.~de-la Cruz-Moreno and H.~Garc\'\i{}a-Compe\'an, ``{Star-triangle type
  relations from $2d$ $\mathcal{N}=(0,2)$ $USp(2N)$ dualities},''
  \href{http://dx.doi.org/10.1007/JHEP01(2021)023}{{\em JHEP} {\bfseries 01}
  (2021) 023}, \href{http://arxiv.org/abs/2008.02419}{{\ttfamily
  arXiv:2008.02419 [hep-th]}}.

\bibitem{Mullahasanoglu:2021xyf}
M.~Mullahasanoglu and N.~Tas, ``{Lens Partition Functions and Integrability
  Properties},'' \href{http://arxiv.org/abs/2112.15161}{{\ttfamily
  arXiv:2112.15161 [hep-th]}}.

\bibitem{Gahramanov:2022qge}
I.~Gahramanov, ``{Integrability from supersymmetric duality: a short review},''
  \href{http://arxiv.org/abs/2201.00351}{{\ttfamily arXiv:2201.00351
  [hep-th]}}.

\bibitem{Bazhanov:2010kz}
V.~V. Bazhanov and S.~M. Sergeev, ``{A Master solution of the quantum
  Yang-Baxter equation and classical discrete integrable equations},''
  \href{http://dx.doi.org/10.4310/ATMP.2012.v16.n1.a3}{{\em Adv. Theor. Math.
  Phys.} {\bfseries 16} (2012) 65--95},
\href{http://arxiv.org/abs/1006.0651}{{\ttfamily arXiv:1006.0651 [math-ph]}}.
%%CITATION = ARXIV:1006.0651;%%.

\bibitem{Kels:2013ola}
A.~P. Kels, ``{A new solution of the star-triangle relation},''
  \href{http://dx.doi.org/10.1088/1751-8113/47/5/055203}{{\em J. Phys.}
  {\bfseries A47} (2014) 055203},
\href{http://arxiv.org/abs/1302.3025}{{\ttfamily arXiv:1302.3025 [math-ph]}}.
%%CITATION = ARXIV:1302.3025;%%.

\bibitem{Kels:2015bda}
A.~P. Kels, ``{New solutions of the star–triangle relation with discrete and
  continuous spin variables},''
  \href{http://dx.doi.org/10.1088/1751-8113/48/43/435201}{{\em J. Phys.}
  {\bfseries A48} no.~43, (2015) 435201},
\href{http://arxiv.org/abs/1504.07074}{{\ttfamily arXiv:1504.07074 [math-ph]}}.
%%CITATION = ARXIV:1504.07074;%%.

\bibitem{Kashaev:2015nya}
R.~Kashaev, ``{The Yang\textendash{}Baxter relation and gauge invariance},''
  \href{http://dx.doi.org/10.1088/1751-8113/49/16/164001}{{\em J. Phys. A}
  {\bfseries 49} no.~16, (2016) 164001},
  \href{http://arxiv.org/abs/1510.03043}{{\ttfamily arXiv:1510.03043
  [math-ph]}}.

\bibitem{Andersen:2014aoa}
J.~E. Andersen and R.~Kashaev, ``{Complex Quantum Chern-Simons},''
  \href{http://arxiv.org/abs/1409.1208}{{\ttfamily arXiv:1409.1208 [math.QA]}}.

\bibitem{Gahramanov:2017idz}
I.~Gahramanov and S.~Jafarzade, ``{Comments on the multi-spin solution to the
  Yang-Baxter equation and basic hypergeometric sum/integral identity},''
  \href{http://dx.doi.org/10.1142/S0217732319501402}{{\em {Mod. Phys. Lett.}}
  {\bfseries A34,no.18,1950140} (2019) },
\href{http://arxiv.org/abs/1710.09106}{{\ttfamily arXiv:1710.09106 [math-ph]}}.
%%CITATION = ARXIV:1710.09106;%%.

\bibitem{Eren:2019ibl}
E.~Eren, I.~Gahramanov, S.~Jafarzade, and G.~Mogol, ``{Gamma function solutions
  to the star-triangle equation},''
  \href{http://dx.doi.org/10.1016/j.nuclphysb.2020.115283}{{\em Nucl. Phys. B}
  {\bfseries 963} (2021) 115283},
  \href{http://arxiv.org/abs/1912.12271}{{\ttfamily arXiv:1912.12271
  [math-ph]}}.

\bibitem{Baxter:1982zz}
R.~J. Baxter, {\em Exactly {S}olved {M}odels in {S}tatistical {M}echanics}.
\newblock Academic, London,
1982.
\newblock
%%CITATION = INSPIRE-1120339;%%.

\bibitem{Bazhanov:2007mh}
V.~V. Bazhanov, V.~V. Mangazeev, and S.~M. Sergeev, ``{Faddeev-Volkov solution
  of the Yang-Baxter equation and discrete conformal symmetry},''
  \href{http://dx.doi.org/10.1016/j.nuclphysb.2007.05.013}{{\em Nucl. Phys.}
  {\bfseries B784} (2007) 234--258},
\href{http://arxiv.org/abs/hep-th/0703041}{{\ttfamily arXiv:hep-th/0703041
  [hep-th]}}.
%%CITATION = HEP-TH/0703041;%%.

\bibitem{Bazhanov:2007vg}
V.~V. Bazhanov, V.~V. Mangazeev, and S.~M. Sergeev, ``{Exact solution of the
  Faddeev-Volkov model},''
  \href{http://dx.doi.org/10.1016/j.physleta.2007.10.053}{{\em Phys. Lett.}
  {\bfseries A372} (2008) 1547--1550},
\href{http://arxiv.org/abs/0706.3077}{{\ttfamily arXiv:0706.3077
  [cond-mat.stat-mech]}}.
%%CITATION = ARXIV:0706.3077;%%.

\bibitem{Dolan:2011rp}
F.~A.~H. Dolan, V.~P. Spiridonov, and G.~S. Vartanov, ``{From 4d superconformal
  indices to 3d partition functions},''
  \href{http://dx.doi.org/10.1016/j.physletb.2011.09.007}{{\em Phys. Lett.}
  {\bfseries B704} (2011) 234--241},
\href{http://arxiv.org/abs/1104.1787}{{\ttfamily arXiv:1104.1787 [hep-th]}}.
%%CITATION = ARXIV:1104.1787;%%.

\bibitem{Spiridonov:2014cxa}
V.~P. Spiridonov and G.~S. Vartanov, ``{Vanishing superconformal indices and
  the chiral symmetry breaking},''
  \href{http://dx.doi.org/10.1007/JHEP06(2014)062}{{\em JHEP} {\bfseries 06}
  (2014) 062},
\href{http://arxiv.org/abs/1402.2312}{{\ttfamily arXiv:1402.2312 [hep-th]}}.
%%CITATION = ARXIV:1402.2312;%%.

\bibitem{Gahramanov:gka}
I.~B. Gahramanov and G.~S. Vartanov,
  \href{http://dx.doi.org/10.1142/9789814449243_0076}{``{Superconformal indices
  and partition functions for supersymmetric field theories},''} in {\em
  {XVIIth Intern. Cong. Math. Phys. 695-703 (2013)}}.
\newblock 2013.
\newblock
\href{http://arxiv.org/abs/1310.8507}{{\ttfamily arXiv:1310.8507 [hep-th]}}.
\newblock
%%CITATION = ARXIV:1310.8507;%%.

\bibitem{Gahramanov:2013xsa}
I.~Gahramanov and G.~Vartanov, ``{Extended global symmetries for 4D $N$ = 1
  SQCD theories},''
  \href{http://dx.doi.org/10.1088/1751-8113/46/28/285403}{{\em J. Phys.}
  {\bfseries A46} (2013) 285403},
\href{http://arxiv.org/abs/1303.1443}{{\ttfamily arXiv:1303.1443 [hep-th]}}.
%%CITATION = ARXIV:1303.1443;%%.

\bibitem{Gahramanov:2015tta}
I.~Gahramanov, ``{Mathematical structures behind supersymmetric dualities},''
  \href{http://dx.doi.org/10.5817/AM2015-5-273}{{\em Archivum Math.} {\bfseries
  51} (2015) 273--286},
\href{http://arxiv.org/abs/1505.05656}{{\ttfamily arXiv:1505.05656 [math-ph]}}.
%%CITATION = ARXIV:1505.05656;%%.

\bibitem{Bozkurt:2018xno}
D.~N. Bozkurt and I.~Gahramanov, ``{Pentagon identities arising in
  supersymmetric gauge theory computations},''
  \href{http://dx.doi.org/10.1134/S0040577919020028}{{\em Teor. Mat. Fiz.}
  {\bfseries 198} no.~2, (2019) 215--224},
\href{http://arxiv.org/abs/1803.00855}{{\ttfamily arXiv:1803.00855 [math-ph]}}.
%%CITATION = ARXIV:1803.00855;%%.

\bibitem{2018}
G.~Sarkissian and V.~P. Spiridonov, ``From rarefied elliptic beta integral to
  parafermionic star-triangle relation,''
  \href{http://dx.doi.org/10.1007/jhep10(2018)097}{{\em Journal of High Energy
  Physics} {\bfseries 2018} no.~10, (Oct, 2018) }.
  \url{http://dx.doi.org/10.1007/JHEP10(2018)097}.

\bibitem{Imamura:2012rq}
Y.~Imamura and D.~Yokoyama, ``{$S^3/Z_n$ partition function and dualities},''
  \href{http://dx.doi.org/10.1007/JHEP11(2012)122}{{\em JHEP} {\bfseries 11}
  (2012) 122},
\href{http://arxiv.org/abs/1208.1404}{{\ttfamily arXiv:1208.1404 [hep-th]}}.
%%CITATION = ARXIV:1208.1404;%%.

\bibitem{Nieri:2015yia}
F.~Nieri and S.~Pasquetti, ``{Factorisation and holomorphic blocks in 4d},''
  \href{http://dx.doi.org/10.1007/JHEP11(2015)155}{{\em JHEP} {\bfseries 11}
  (2015) 155},
\href{http://arxiv.org/abs/1507.00261}{{\ttfamily arXiv:1507.00261 [hep-th]}}.
%%CITATION = ARXIV:1507.00261;%%.

\end{thebibliography}\endgroup

\documentclass[11pt,a4paper]{article}
\usepackage{jheppub}
\usepackage{amsmath}
\usepackage[most]{tcolorbox}
\usepackage{dsfont}
\usepackage{ulem}
\usepackage{natbib}
\usepackage{xcolor}
\usepackage[hang,flushmargin]{footmisc}
\usepackage{tikz-cd}
\usepackage{enumitem}
\usepackage{amsfonts,amssymb, amscd,amsmath,latexsym,amsbsy,bm}
\usepackage{stmaryrd}
\usepackage{todonotes}
\usepackage{stackrel}
\usepackage{amsthm}

\usepackage{graphicx}
\newcommand{\imi}{\mathsf{i}}

\title{A comment on the solutions of the generalized Faddeev-Volkov model}

\author{Mehmet Dede}
\affiliation{
	 {Department of Physics, Bogazici University, 34342 Bebek, Istanbul, Turkey}\\[-0.5cm]

}

\emailAdd{mehmet.dede@boun.edu.tr}

\abstract{We consider two recent generalizations of the Faddeev-Volkov model, which is exactly solvable Ising-type lattice spin model. The first generalization based on using of the non-compact quantum dilogarithm over Pontryagin self-dual LCA group $\mathbb{R}\times \mathbb{Z}/N\mathbb{Z}$, and  another one constructed in a recent study via the gauge/YBE correspondence. We show that weight functions of these models obtained by different techniques are the same up to a constant.}

\keywords{generalized Faddeev-Volkov model, star-triangle relation}

\begin{document}
	\maketitle
	\flushbottom

\section{Introduction}
In recent years, there have been several progress in constructing new integrable systems in terms of Yang-Baxter equations and generalizing known solutions. Some studies employ the gauge/YBE correspondence to connect supersymmetric gauge theories and integrable lattice models of statistical mechanics, e.g. \cite{,Spiridonov:2010em,Yamazaki:2012cp,Gahramanov:2015cva,Gahramanov:2016ilb,Gahramanov:2017ysd,Kels:2017vbc,Yamazaki:2018xbx,Bozkurt:2020gyy,de-la-Cruz-Moreno:2020xop,Mullahasanoglu:2021xyf,Gahramanov:2022qge}, whereas others use special functions and algebraic techniques, see, e.g. \cite{Bazhanov:2010kz,Kels:2013ola,Kels:2015bda,Kashaev:2015nya,Andersen:2014aoa,Gahramanov:2017idz}. It turns out that, under the particular case, solutions derived from using two different constructions might take similar forms \cite{Eren:2019ibl}.

In the work \cite{Kashaev:2015nya}, Kashaev generalized the solution of the Faddeev-Volkov statistical model, utilizing the quantum dilogarithm $\phi(x,m)$ over LCA group $A$, and the Boltzmann weight function
\begin{equation}\label{weight}
     W(x,y)=\frac{\phi(x-y)}{\phi(x+y)}\langle x;y \rangle = \frac{\phi(x-y) \phi(-x-y)}{\phi(0)^2 \langle x \rangle \langle  y\rangle} \; ,
\end{equation}
where $x$ and $y$ take values in specified self-dual LCA group $A$, and $\langle \cdot \rangle : A \rightarrow \mathbb{T}$ is Gaussian exponential, whose existence is implied by the definition of self-duality of LCA group \cite{Kashaev:2015nya}. The weight function (\ref{weight}) solves the operator Yang-Baxter equation
\begin{equation}
    W(\textbf{p},x) W(\textbf{q},x+y)W(\textbf{p},y)=W(\textbf{q},y) W(\textbf{p},x+y)W(\textbf{q},x)\; ,
\end{equation}
where \textbf{p} and \textbf{q} are self-adjoint operators in Hilbert space and they act as differentiation and multiplication operators respectively. Therefore, one can construct the corresponding integrable lattice spin model of statistical mechanics \cite{Baxter:1982zz}. 

We set our framework around a particular case $A=\mathbb{R}\times \mathbb{Z}/N\mathbb{Z}$. Over LCA group $\mathbb{R}\times \mathbb{Z}/N\mathbb{Z}$ the Gaussian exponential takes the form $\langle (x,m) \rangle=e^{-\pi \imi x^2 }e^{-\pi \imi m(m+N)/N}$ and explicit version of quantum dilogirthm $\phi(x,m)$  can be written as \cite{Andersen:2014aoa,Kashaev:2015nya}:
\begin{equation}\label{quantumdilog}
\phi(x,m)=\prod_{j=0}^{N -1}\Phi_b\left(\frac{x}{\sqrt{N }}+\imi(1-N^{-1})\left(\frac{b+b^{-1}}{2}\right)-\imi b^{-1}\frac{j}{N }-\imi b\left\{\frac{j+m}{N }\right\}\right) \;,
\end{equation}
where $\{x\}$ represents the fractional part of $x$. 

On the other hand, recently, in \cite{Bozkurt:2020gyy} there was obtained a new solution to the star-triangle equation (a special form of the Yang-Baxter equation), which is a generalization of the Faddeev-Volkov model \cite{Bazhanov:2007mh,Bazhanov:2007vg} using gauge/YBE correspondence. Namely in order to construct a new solution to the star-triangle equation, authors of \cite{Bozkurt:2020gyy} used three-dimensional supersymmetric duality for $\mathcal N=2$ theories on lens space $S^3/\mathbb{Z}_r$.  

In this work, we show that two generalizations of the Faddeev-Volkov model constructed in \cite{Kashaev:2015nya} and \cite{Bozkurt:2020gyy} are identical in regard to their integrability properties.

\section{Hyperbolic gamma function and its properties}

%First, we will define some of the special functions that will be of use and provide the relations between them. 
To define these special functions, we need to introduce the q-Pochhammer symbol
\begin{align}
    (z,q)_\infty=\prod_{i=1}^\infty (1-z q^i), &&  \vert q\vert< 1 \;, && z \in \mathbb{C}\;.
\end{align}
Defining the q-Pochhammer symbol, we can write an explicit version of the non-compact quantum dilogartihm, which is the function Kashaev uses in his work \cite{Kashaev:2015nya} while generalizing the Faddeev-Volkov's solution \cite{Bazhanov:2007mh, Bazhanov:2007vg},
\begin{align}
    \Phi_b(z)= \frac{(-q_b e^{2\pi \imi b x};q^2_b)_\infty}{(-\Bar{q}_b e^{2 \pi \imi b^{-1} x};\Bar{q}_b^2)_\infty},  && q_b=e^{\pi \imi b^2 }, && \Bar{q}_b=e^{-\pi \imi b^{-2}}, && Im(b^2)>0 \;.
\end{align}

There is a similar function called the hyperbolic gamma function\footnote{This function appears in several areas of mathematical physics, this form of notation is often found in works of supersymmetry community, see, e.g. \cite{Dolan:2011rp,Spiridonov:2014cxa,Gahramanov:gka, Gahramanov:2013xsa, Gahramanov:2015tta,Bozkurt:2018xno}.} $\gamma^{(2)}(z,b,b^{-1})$. Using the q-Pochhammer symbol
\begin{align}
    	\gamma^{(2)}(z;b,b^{-1})=e^{-\frac{\pi i}{2}B_{2,2}(z;b,b^{-1})}\frac{(e^{2\pi i b}q^2_b; q^2_b)}{(e^{2\pi i\frac{z}{b}};\Bar{q}_b^2)}  \;,
\end{align}
where $B_{2,2}$ is the second diagonal Bernoulli polynomial,
\begin{equation}
    B_{2,2}(z;b,b^{-1})=z^2-z(b+b^{-1})+\frac{b^2+3+ b^{-2}}{6 }\;.
\end{equation}
The relation between hyperbolic gamma function $\gamma^{(2)}(z,b,b^{-1})$ and non-compact quantum dilogarithm $\Phi_b(z)$ can be stated as
\begin{equation}\label{eqeq}
    \Phi_b(z)=e^{-\pi \imi B_{2,2}(\frac{1}{2}(b+b^{-1})+iz,b,b^{-1})}\gamma^{(2)}(\frac{1}{2}(b+b^{-1})+iz;b,b^{-1})\;.
\end{equation}

One can also express the hyperbolic gamma function in integral form
\begin{equation}
   \gamma^{(2)}(z;b,b^{-1})=\exp{\left(-\int_{0}^{\infty}\frac{dx}{x}\left[\frac{\sinh{x(2z-b-b^{-1})}}{2\sinh{(xb)}\sinh{(xb^{-1})}}-\frac{2z-b-b^{-1}}{2x}\right]\right)} \;,
\end{equation}
where $Re(b),Re(b^{-1})>0 $ and $Re(b+b^{-1})> Re(z)>0$.

The hyperbolic gamma function has several practical properties, however, we will only use two of them:
\begin{align}\label{inv}
\gamma^{(2)}(z;b,b^{-1})\gamma^{(2)}(b+b^{-1}-z;b,b^{-1})=1 \\
	\label{scaling}
\gamma^{(2)}(z;b,b^{-1})=\gamma^{(2)}(u z;u b,u b^{-1})\;.
\end{align}

\section{Star-triangle relation}

In this section, we show that Kashaev's weight function differs from the Boltzmann weights constructed in \cite{Bozkurt:2020gyy} only by a scaling factor. To this end, we express $\Phi_b(z)$ in terms of the hyperbolic gamma function (\ref{eqeq}) and exploit the mentioned properties of the gamma function.

First, we replace the non-compact quantum dilogarithm $\Phi_b(z)$ with the hyperbolic gamma functions in the expression (\ref{quantumdilog})
\begin{align} \nonumber
    \phi(x,m)=\exp{\left(-\pi \imi \sum\limits_{j=0}^{N-1} B_{2,2}\left(\frac{\imi x}{\sqrt{N}}+\frac{(b+b^{-1})}{2}N^{-1}+\frac{b^{-1}j}{N}+b \left\{\frac{j+m}{N }\right\};b,b^{-1}\right)\right)}\\
    \prod_{j=0}^{N-1}\gamma^{(2)}\left(\frac{\imi x}{\sqrt{N}}+\frac{(b+b^{-1})}{2}N^{-1}+\frac{b^{-1}j}{N}+b \left\{\frac{j+m}{N }\right\};b,b^{-1}\right)\;.
\end{align}

In order to manipulate and use the properties of $\gamma^{(2)}(z,b,b^{-1})$, one may want to eliminate the fractional part in the argument. Thus, considering the fractional part, it can easily be seen that the product can be rewritten as two separated parts (for the simplicity let us denote $ \frac{\imi x}{\sqrt{N}}+\frac{(b+b^{-1})}{2}N^{-1}$  as $ \frac{y}{N}$)
\begin{align} \nonumber 
    \prod_{j=0}^{N-1}\gamma^{(2)}\left(\frac{y}{N}+\frac{b^{-1}j}{N}+b \left\{\frac{j+m}{N }\right\};b,b^{-1}\right)=\prod_{j=0}^{m-1}\gamma^{(2)}
\left({\frac{y}{N} }+b\left(1-{\frac{m}{N}}\right)+(b+b^{-1}){\frac{j}{N}};b,b^{-1}\right)\\ \label{productN}
\times\prod_{j=0}^{N-m-1}\gamma^{(2)}
\left({\frac{y}{N}}+{\frac{m}{N}}b+(b+b^{-1}){\frac{j}{N}};b,b^{-1}\right)\ \;.
\end{align}

Considering the product of two standard hyperbolic gamma functions in different bases, and using the limiting behavior mentioned in \cite{2018}, it can be shown that the expression on the right-hand side (\ref{productN}) is indeed equal to the product of two hyperbolic gamma functions. For the derivation of this relation and further relations, see \cite{Gahramanov:2016ilb,2018,Imamura:2012rq,Nieri:2015yia} 
\begin{align}\nonumber
    \prod_{j=0}^{N-1}\gamma^{(2)}\left(\frac{y}{N}+\frac{b^{-1}j}{N}+b \left\{\frac{j+m}{N }\right\};b,b^{-1}\right)=\gamma^{(2)}\left({y+m \frac{b}{N}};b,{\frac{b+b^{-1}}{N}}\right)\\
\gamma^{(2)}\left({\frac{y+(N-m)b^{-1}}{N}};b^{-1},{\frac{b+b^{-1}}{N}}\right)\;.
\end{align}

Therefore, the quantum dilogarithm $\phi(x,m)$ can be written as
\begin{align}\nonumber 
    \phi(x,m)=\exp{\left(-\pi \imi \sum\limits_{j=0}^{N-1} B_{2,2}\left(\frac{\imi x}{\sqrt{N}}+\frac{(b+b^{-1})}{2}N^{-1}+\frac{b^{-1}j}{N}+b \left\{\frac{j+m}{N }\right\};b,b^{-1}\right)\right)}
    \\\gamma^{(2)}\left(\frac{\imi x}{\sqrt{N}}+\frac{(b+b^{-1})}{2}N^{-1}+{\frac{m b}{N}};b,{\frac{b+b^{-1}}{N}}\right)\nonumber
    \\\label{summ}
 \gamma^{(2)}\left(\frac{\imi x}{\sqrt{N}}+\frac{(b+b^{-1})}{2}N^{-1}+{\frac{(N-m)b^{-1}}{N}};b^{-1},{\frac{b+b^{-1}}{N}}\right) \;.
\end{align}
Calculating the summation is nearly an identical task with calculating the product. One should separate the summation into two parts from $j=0$ to $m-1$ and from $0$ to $N-m-1$. Rest of the calculation follows from the definition of second diagonal Bernoulli polynomial $B_{2,2}(z, b,b^{-1})$ and elementary summation identities. After computing the summation, we insert the spin values which take values in LCA group. In Kashaev's setting, there are two different inputs of the weight function (\ref{weight}). The variable $x$ stands for differences of discrete spin ($m_1$ and $m_2$) and continuous spin ($x_1$ and $x_2$), hence $x=(x_1-x_2,m_1-m_2)$. The variable $y$ denotes the rapidity lines $(-\alpha,0)$.

Inserting these variables, we use scaling property of the hyperbolic gamma function (\ref{scaling}) where $u= -\imi N $. Finally, after the following change of variables
\begin{gather}
    x_1-x_2-\alpha \rightarrow -\imi \frac{x_1-x_2-\alpha}{\sqrt{N}}+\imi \frac{b+b^{-1}}{\sqrt{N}} \;, \\
    -x_1+x_2-\alpha \rightarrow -\imi \frac{x_1+x_2-\alpha}{\sqrt{N}}+\imi \frac{b+b^{-1}}{\sqrt{N}}\;,
\end{gather}
we observe that the weight function in Kashaev's settings is indeed a scaled version of the Boltzmann weight constructed in \cite{Bozkurt:2020gyy}
\begin{align}\nonumber
    W(x,y)=W_\alpha (x_1,x_2,m_1,m_2)=C e^{-\pi \imi (m_1 +m_2)}\gamma^{(2)}\left(-\imi (-\alpha +x_1-x_2)-\imi b(m_1-m_2)  ; -\imi b N , -\imi(b+b^{-1})\right)\times\\\nonumber
\gamma^{(2)}\left(-\imi (-\alpha+x_1-x_2)-\imi b^{-1} (N-(m_1-m_2));-\imi N b^{-1},-\imi(b+b^{-1}) \right)\times\\ \nonumber \gamma^{(2)}\left(-\imi (-\alpha-x_1+x_2)-\imi b(-m_1+m_2)  ; -\imi b N , -\imi (b+b^{-1})\right)\times\\\nonumber
\gamma^{(2)}\left(-\imi (-\alpha-x_1+x_2)-\imi b^{-1} (N-(-m_1+m_2)) ;-\imi N b^{-1},-\imi (b+b^{-1}) \right) \;,
\end{align}
with $C=\exp{\left[\frac{i \pi  b^2}{24 N}+\frac{i \pi }{24 b^2 N}-\frac{i \pi  N}{12}+\frac{i \pi }{12 N}\right] }$. Even though two Boltzmann weights differ from each other by a constant factor, they possess identical integrability properties as they satisfy the same star triangle equation 
\begin{align*}
    \frac{1}{\imi N}\sum\limits^{\lfloor N/2 \rfloor}_{u=-\lfloor N/2 \rfloor} e^{-2 \pi \imi u } \int\limits^\infty_{-\infty} dz W_{\alpha_1}(x_1,z, m_1,u) W_{\alpha_2}(x_1,z, m_1,u) W_{\alpha_3}(x_1,z, m_1,u)\\
    =W_{\alpha_1+\alpha_2}(x_1,x_2, m_1,m_2)W_{\alpha_1+\alpha_3}(x_1,x_2, m_1,m_2)W_{\alpha_2+\alpha_3}(x_1,x_2 ,m_1,m_2) \;.
\end{align*}

There are many solutions to the star-triangle relation obtained via the gauge/YBE correspondence, it would be interesting to study them from the algebraic point of view as in \cite{Kashaev:2015nya}.   

\bibliographystyle{utphys}
\bibliography{pentagon}

\end{document}